OpenStudy (anonymous):

A rectangular pen for a pet is 3ft longer that it is wide, give possible values for the width W of the pen if its area must be between 180 and 460 sqaure feet, inclusivley.

OpenStudy (anonymous):

If area is 180 W * ( W + 3) = 180 w^2 + 3w = 180 w = 12 if area is 460, W^2 + 3w = 460, w = 20 12<W<20

OpenStudy (anonymous):

so that would mean 12ft and 20 ft

OpenStudy (anonymous):

No, the width can lie ANYWHERE between 12ft and 20ft. He tested the extreme cases here, and reasoned that a value of W somewhere between 12ft and 20ft would result in an area greater than 180 but less than 460. If you'd like to confirm this fact, try graphing\[Area=w*(w+3)=w^2+3w\] Look at what values of w output values that are between 180 and 460 and you should see that 12ft and 20ft indeed bound these values! I hope that helps solidify your understanding!